Can someone out there help me solve this problem?
Is the rate 1.2 inches of rain/5 hours equivalent to the rate 2.1 inches of rain/8 hours?
Thanks,
Lyn

Answer by tutorcecilia(2144)

No.
If you divide each fraction you can than compare the amount of rain that falls in only one hour:
.
1.2 inches of rain/5 hours = 1.2/5 = .2400 inches/1 hour (the least rain fall)
.
2.1 inches of rain/8 hours = 2.1/8 =.2625 inches/1 hour (the most rain fall)